OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 39 miles at a certain speed. The canoeist travels 19 miles on the second part of the trip at a speed 5 mph slower. The total time from the trip is 3 hrs. What was the speed on each part of the trip

OpenStudy (anonymous):

at=39 (a-5)s =19 t+s=3

OpenStudy (anonymous):

imranmeah91 that was not correct

OpenStudy (dumbcow):

haha those are the equations to use, not the answer distance = rate*time

OpenStudy (dumbcow):

a is rate or speed s and t are times for each trip respectively

OpenStudy (dumbcow):

solve for time for both trips, then set sum equal to 3 at=39 --> t = 39/a (a-5)s = 19 --> s = 19/(a-5) t+s = 3 \[\rightarrow \frac{39}{a}+\frac{19}{a-5} = 3\] combine fractions \[\frac{39(a-5) + 19a}{a(a-5)} = 3\] \[39(a-5) +19a = 3a(a-5)\] distribute and add like terms \[39a -195 +19a = 3a^{2} -15a\] \[3a^{2}-73a+195 = 0\] Solve using quadratic formula